OpenStudy (90890908967777878888):

order the numbers from least to greatest 3/8 19/24 2/3 ill give u an medal

OpenStudy (anonymous):

3/8 = 0.375, 19/24 = 0.791, 2/3 = 0.667

OpenStudy (anonymous):

So it would be 3/8, 2/3, and then 19/24

OpenStudy (90890908967777878888):

thx